# {textblock*}

### v2

@@ I Introduction
At the emission of the cosmic microwave background (CMB), four fluids contribute to the energy of the Universe: photons, baryons, neutrinos and dark matter. Taking a bottom-up perspective, it is therefore natural to assume that there are four independent initial conditions for perturbations to the homogeneous and isotropic background cosmology. In terms of gauge invariant quantities, they may for example be given by a curvature perturbation $\zeta$ and isocurvature perturbations between the photons and the other fluids, $S_{\gamma X}$ with $X\in\{\nu,\mathrm{DM},\mathrm{b}\}$[1].
@@
To this day, all observations of anisotropies are in agreement with the existence of only a non-vanishing initial curvature perturbation $\zeta\neq 0\,,S_{\gamma X}=0$ In terms of cosmological histories this is easily motivated. In any scenario where there is only one single degree of freedom during inflation or at some point after inflation, i.e., the whole energy content thermalizes, this so called adiabatic initial condition is realized [2, 3].
From a top-down perspective, constraining isocurvature therefore corresponds to testing these most simple cosmologies, while the discovery of isocurvature would point to a much richer cosmology, since throughout its evolution there must be multiple non-thermalized sectors present. With this motivation in mind, one may try to search for the four independent initial conditions mentioned above. This corresponds to generalizing the curvature power spectrum $\mathcal{P}_{\zeta}(k)$ which quantizes the correlator $\langle\zeta(k)\zeta(k)\rangle$ into a symmetric 4x4 matrix that further gives the correlators $\langle\zeta(k)S_{\gamma X}(k)\rangle$ and $\langle S_{\gamma X}(k)S_{\gamma Y}(k)\rangle$[4]. To our knowledge such a search has never been carried out though and we expect that this is due to the large number of additional parameters one has to infer from cosmological data. While $\mathcal{P}_{\zeta}(k)$ is commonly given in terms of a normalization and a tilt, so two free parameters, doing the same with all other entries of this symmetric matrix would result in 18 additional free parameters.
@@
In most previous searches people have instead considered generalizations where besides $\zeta\neq 0$, one additional linear combination of the $S_{\gamma X}$ is non-zero  [^1] WMAP data was analyzed in terms of up to two isocurvature initial conditions with arbitrary correlations [49, 50, 51].. With this restriction, the power spectrum is generalized to a 2x2 matrix and with the same counting as above one introduces four additional parameters only. From the top-down one may try to justify this procedure in the following way: So far only curvature perturbations have been observed and therefore it seems plausible that there is a hierarchy between the different perturbations. If this is the case going from a cosmological history where at early times only one degree of freedom was of relevance, as a first step, it is enough to consider one where there are two independent sectors at all times. The immediate questions that come up in this context are which cosmological history leads to which linear combination of $S_{\gamma X}$ and whether any of these combinations are particularly well motivated.
In Section II we consider some cosmological histories of increasing complexity and discuss the isocurvature mode generated by them. We find that cosmologies with neutrino isocurvature generically feature fully correlated admixtures of matter isocurvature. This motivates us to conduct a search for this generalized isocurvature using CMB and large scale structure (LSS) data, where $S_{\gamma\nu}/S_{\gamma\mathrm{DM}}$ is left as a free parameter, in Section III .
@@ II Overview on cosmic histories
Let us now return to the question whether any linear combinations of isocurvature are particularly well motivated. First note that the so called compensated matter perturbation, which is defined by the isocurvature of the total matter content with respect to the total radiation vanishing and corresponds to
@@
is not observable at linear order in the CMB [6, 7]. It is therefore only constrained at amplitudes much larger than the observed adiabatic ones [8, 9, 10]. In this paper we restrict ourselves to magnitudes of isocurvature smaller than the observed CMB fluctuations, $S\lesssim 10^{-5}$. This allows us to recast any fluctuation of baryons and dark matter into an effectively unconstrained compensated matter perturbation and an overall matter fluctuation that may be expressed by a dark matter fluctuation, while choosing $S_{\gamma b}=0$ The question of what isocurvature perturbation to consider is then reduced to a choice of the ratio $S_{\gamma\nu}/S_{\gamma\mathrm{DM}}$
@@
Previous works [11, 12, 13, 8, 14, 15] have only considered specific choices of the angle $\varphi$ The initial condition called dark matter isocurvature corresponds to $\varphi=0$ with our definitions. The one referred to as neutrino isocurvature is however arranged such that there is no isocurvature between the total radiation and matter. This corresponds to the somewhat odd combination $\varphi~{}=~{}\arctan[(1+\rho_{\gamma}/\rho_{\nu})/(1+\rho_{b}/\rho_{\mathrm{DM}})]\approx 1.12$ Further the normalization is chosen such that the isocurvature fraction $f_{\mathrm{iso}}$ needs to be multiplied by a factor of $1.4$ to compare to our convention. We will refer to this choice of $\varphi$ as compensated neutrino isocurvature in analogy to the compensated matter perturbation.
II.1 Neutrino isocurvature from a dark sector
To see whether these values of $\varphi$ are particularly well motivated, we compare cosmologies of increasing complexity in Fig. 1. The first one, (a), is the standard single-field inflation scenario with inflaton $\Phi\,$. In this model, at early times there can only be a time difference between the equal energy slicing of the inflaton sector with respect to spatially flat slices, which leads to the fluctuations being adiabatic. The fluctuations then stay adiabatic throughout all of the following evolution. This holds true for much more complicated inflationary sectors as long as they reheat a common fluid at some point [3, 2]. An example might be multiple fields $\Psi_{s}$ driving inflation as illustrated in Fig. 1.
The following three examples all feature two independent, non-thermalized fluids throughout the evolution from inflation until today and therefore feature besides curvature fluctuations one independent linear combination of isocurvature. In example (b), the second sector $\Psi$ later on acts as cold dark matter (DM). A minimal example would be scalar DM produced via the misalignment mechanism [16, 17, 18, 19, 20, 21]. Since the photons and neutrinos are in thermal equilibrium at the time when they decouple, the relative isocurvature between them is $\mathcal{S}_{\gamma\nu}=0\,$. One therefore only has DM isocurvature $\mathcal{S}_{\gamma\mathrm{DM}}\neq 0\,$, as one may have naively expected.
In (c) and (d), we show examples of cosmologies that realize a non-zero neutrino isocurvature. As just discussed, neutrinos are in thermal equilibrium at times well before big bang nucleosynthesis (BBN). In order to achieve fluctuations in the neutrino density one must therefore either consider fluctuations in the lepton number or consider a separate sector that at times after BBN contributes to the neutrinos. We will first discuss the second possibility.
Note that while neutrino masses are expected to become relevant to LSS surveys in the near future [22, 23, 24], to this day neutrinos may be modeled as free streaming radiation, only interacting gravitationally at all times after BBN. Let us therefore consider the simplest case, and instead of neutrinos consider a dark sector that at the time of CMB emission behaves as free streaming, dark radiation. Neutrinos and the dark sector are then indistinguishable to CMB and LSS probes. To this end we define an effective neutrino density and isocurvature perturbation
where we assumed $S_{\gamma\nu}=0$. One may naively think that scenario (c) in the absence of direct interaction (indicated by gray arrows in Fig. 1) corresponds to $\mathcal{S}_{\gamma\nu_{\mathrm{eff}}}\neq 0\,$ and $\mathcal{S}_{\gamma\mathrm{DM}}=0$, while (d) corresponds to $\mathcal{S}_{\mathrm{DM}\,\mathrm{DR}}=0$, which is equivalent to $S_{\gamma\mathrm{DM}}=\rho_{\nu_{\mathrm{eff}}}/\rho_{\mathrm{DR}}\,S_{\gamma\nu_{\mathrm{eff}}}$.
It is obvious that if we allow for direct interactions between the fluids, any ratio of $\mathcal{S}_{\gamma\nu_{\mathrm{eff}}}/\mathcal{S}_{\gamma\mathrm{DM}}$ may be achieved. For example if we introduce in (c) an interaction that partially thermalizes the DM and DR sector, $\mathcal{S}_{\mathrm{DM}\,\mathrm{DR}}$ would be reduced, and depending on the interaction strength, one can interpolate to any value of $\mathcal{S}_{\gamma\nu_{\mathrm{eff}}}/\mathcal{S}_{\gamma\mathrm{DM}}$ between the ones of example (c) and (d).
In an upcoming paper [25], we will further concentrate on example (c) and show that, contrary to the naive expectation, even in the absence of direct interactions, a nonzero $S_{\gamma\mathrm{DM}},\,S_{\gamma\mathrm{b}}$ can be induced. Using the separate universe approach [26, 27, 28, 29] one can roughly explain the effect as follows: In the presence of dark radiation isocurvature on super horizon scales, the Universe may be viewed as a collection of patches with varying dark radiation density at a given density in the standard model sector (the one originating from $\Phi$ in Fig. 1 (c)). In most models of dark matter production [30] or baryogenesis [31], the yield crucially depends on the ratio between particle physics rates and the Hubble rate. The particle physics rates themselves only depend on the density in the standard model sector (which are usually both given as a function of temperature) and are therefore the same for all patches with the above slicing. The Hubble rate, on the other hand, depends on the energy density in both sectors and therefore varies. This results in varying yields in different patches, a non-vanishing matter isocurvature. Similar effects have been discussed in the context of the curvaton in Refs. [32, 33], and Ref. [14] pointed out that the dark radiation scenario leads to varying element abundances during BBN in the same way.
It should be noted that in the purely gravitationally coupled cases, the amount of induced matter isocurvature is proportional to the relative energy density in the dark sector $\Omega_{\mathrm{DR}}$ at the time of dark matter production and baryogenesis, respectively. The effect might therefore be suppressed if e.g. the dark sector behaves matter-like initially and only later on acts as radiation, resulting in $S_{\gamma\mathrm{DM}}\simeq 0$. This is also the reason why these complications do not appear in case (b) of dark matter isocurvature.
II.2 Neutrino isocurvature from lepton number fluctuations
[…]
A lepton number conserving decay at earlier times only leads to an increase of the temperature in the primordial plasma and would not induce neutrino isocurvature. If the decay induces fluctuations in the lepton number to photon ratio, this may still lead to fluctuations in the neutrino energy density. Since this dependence starts at quadratic order only, a sizable background lepton asymmetry is required in order for its fluctuations to induce perturbations in the energy density at linear order [32, 12]. The fluctuations in the energy density again lead to the gravitational effects that induce matter fluctuations. Furthermore, if the large lepton number is linked to the generation of the observed baryon asymmetry, one would expect an imprint of fluctuations on the baryon number as well  [^2] In scenarios where the SM plasma reaches temperatures well above the electroweak scale, sphaleron processes equilibrate the lepton and baryon asymmetry, favoring a tiny lepton asymmetry [52]. The generation of the baryon asymmetry would need to be a much weaker residual effect.. We therefore reach the same conclusion that neutrino isocurvature is in general accompanied by matter isocurvature.
@@ III.1 Computation of cosmological observables
In order to find the effect of isocurvature on the CMB and other cosmological observables, one needs to solve the evolution equations of the perturbations as the relevant scales enter the horizon. This is commonly done in a fixed gauge. The isocurvature mode in question then corresponds to a set of initial conditions that one has to determine in the gauge of choice. Further, since the evolution equations typically feature singularities as $\tau\rightarrow 0$, one has to analytically expand the solution to finite conformal time $\tau$ before the equations can be solved numerically. In Appendix A we give this expansion in synchronous gauge for $S_{\gamma\mathrm{DM}}=1,\,S_{\gamma\nu}=0$ and $S_{\gamma\mathrm{DM}}=0,\,S_{\gamma\nu}=1$ The general case may then be obtained by adding the solutions with weights $\cos(\varphi)$ and $\sin(\varphi)$ , respectively. This approach is straightforward to generalize to more exotic scenarios containing e.g. self-interacting instead of free-streaming dark radiation [15].
We use a modified version of the code CLASS [35] to solve the evolution equations of the perturbations and compute all relevant CMB and LSS observables. In Fig. 2 we show the spectrum of lensed temperature correlations in the CMB for a fiducial $\mathrm{\Lambda CDM}$ cosmology using the best fit parameters from Planck [36], but for simplicity taking a flat primordial spectrum, i.e., setting the scalar index $n=1$ Besides the adiabatic initial conditions that are dominating our Universe, we show four different scenarios with isocurvature instead of the adiabatic mode at the same amplitude.
@@
It becomes apparent that the resulting spectra differ in the position of the acoustic peaks as well as their heights. It is therefore clear that CMB data is able to distinguish between different values of $\varphi$ In the event of detection, $\varphi$ can therefore be determined and would serve as an invaluable piece of information when trying to disentangle the various cosmologies discussed above. On the other hand, since the signatures differ quiet drastically, it is easy to imagine that an isocurvature signal may be missed when narrowing the search to one $\varphi$ only. The same conclusion is reached in Appendix B, where we consider CMB polarization and LSS.

### v3

# Generalized neutrino isocurvature

@@ I Introduction
At the emission of the cosmic microwave background (CMB), four fluids contribute to the energy of the Universe: photons, baryons, neutrinos and dark matter. Taking a bottom-up perspective, it is therefore natural to assume that there are four independent initial conditions for perturbations to the homogeneous and isotropic background cosmology. In terms of gauge invariant quantities, they may for example be given by a curvature perturbation $\zeta$ and isocurvature perturbations between the photons and the other fluids [11],
$S_{\gamma X}=3(\zeta_{\gamma}-\zeta_{X})=-3H\left(\frac{\delta\rho_{\gamma}}{\dot{\rho}_{\gamma}}-\frac{\delta\rho_{X}}{\dot{\rho}_{X}}\right)\,$, (1)
where $X\in\{\nu,\mathrm{DM},\mathrm{b}\}$, $H$ is the Hubble parameter and $\delta\rho_{\gamma,X}$ the energy density perturbation of the background energy density $\rho_{\gamma,X}$.
@@
To this day, all observations of anisotropies are in agreement with the existence of only a non-vanishing initial curvature perturbation $\zeta\neq 0\,,S_{\gamma X}=0$ . In terms of cosmological histories this is easily motivated. In any scenario where there is only one single degree of freedom during inflation or at some point after inflation, i.e., the whole energy content thermalizes, this so called adiabatic initial condition is realized [53, 52].
@@
From a top-down perspective, constraining isocurvature therefore corresponds to testing these most simple cosmologies, while the discovery of isocurvature would point to a much richer cosmology, since throughout its evolution there must be multiple non-thermalized sectors present. Given that so far only curvature perturbations have been observed, it seems plausible that there is a hierarchy between the different perturbations. If this is the case, going from a cosmological history where at early times only one degree of freedom was of relevance, as a first step, it seems logical to consider one where there are two independent sectors at all times leading to one combination of isocurvature fluctuations besides the adiabatic one. The immediate questions that come up in this context are which cosmological history leads to which linear combination of $S_{\gamma X}$ and whether any of these combinations are particularly well motivated. In an attempt to answer this question, we consider cosmological histories of increasing complexity in Section II (cf. Fig. 1). We find that cosmologies with neutrino isocurvature generically feature fully correlated admixtures of matter isocurvature.
From the bottom-up instead, one may try to keep a search for the four independent initial conditions mentioned above as general as possible. This corresponds to generalizing the curvature power spectrum $\mathcal{P}_{\zeta}(k)$ which quantizes the correlator $\langle\zeta(k)\zeta(k)\rangle$ into a symmetric 4x4 matrix that further gives the correlators $\langle\zeta(k)S_{\gamma X}(k)\rangle$ and $\langle S_{\gamma X}(k)S_{\gamma Y}(k)\rangle$[18]. While $\mathcal{P}_{\zeta}(k)$ is commonly given in terms of a normalization and a tilt, so two free parameters, doing the same with all other entries of this symmetric matrix would result in 18 additional free parameters  [^1] The correlation matrix has to be kept positive definite. See Ref. [2] for details on how to ensure this constraint.. While for WMAP, this most general search has been carried out [19, 17, 47, 13], this is not the case for the best data available, currently coming from Planck [7].
The Planck analysis [2, 3, 7] has instead considered generalizations where besides $\zeta\neq 0$, one additional linear combination of the $S_{\gamma X}$ is non-zero. With this restriction, the power spectrum is generalized to a 2x2 matrix and with the same counting as above one introduces four additional parameters only. To such a search our work on the origin of particular combinations of isocurvature is of immediate importance, since it allows one to interpret the results in terms of specific cosmological histories. The insight from Section II that neutrino isocurvature is generically accompanied by arbitrary amounts of matter isocurvature, motivates us to conduct a search for this generalized isocurvature using CMB and large scale structure (LSS) data, where $S_{\gamma\nu}/S_{\gamma\mathrm{DM}}$ is left as a free parameter, in Section III .
@@ II Overview on cosmic histories
To see whether any specific combinations of isocurvature modes are particularly well motivated, we compare cosmologies of increasing complexity in Fig. 1. The first one, (a), is the standard single-field inflation scenario with inflaton $\Phi\,$. In this model, at early times there can only be a time difference between the equal energy slicing of the inflaton sector with respect to spatially flat slices, which leads to the fluctuations being adiabatic. The fluctuations then stay adiabatic throughout all of the following evolution. This holds true for much more complicated inflationary sectors as long as they reheat a common fluid at some point [52, 53]. An example might be multiple fields $\Psi_{s}$ driving inflation as illustrated in Fig. 1.
The following three examples all feature two independent, non-thermalized fluids throughout the evolution from inflation until today and therefore feature besides curvature fluctuations one independent linear combination of isocurvature. In example (b), the second sector $\Psi$ later on acts as cold dark matter (DM). A minimal example would be scalar DM produced via the misalignment mechanism [48, 1, 24, 40, 32, 20]. Since the photons and neutrinos are in thermal equilibrium at the time when they decouple, the relative isocurvature between them is $S_{\gamma\nu}=0\,$. One therefore only has DM isocurvature $S_{\gamma\mathrm{DM}}\neq 0\,$, as one may have naively expected.
In (c) and (d), we show examples of cosmologies that realize a non-zero neutrino isocurvature. As just discussed, neutrinos are in thermal equilibrium at times well before big bang nucleosynthesis (BBN). In order to achieve fluctuations in the neutrino density one must therefore either consider fluctuations in the lepton number or consider a separate sector that at times after BBN contributes to the neutrinos. We will discuss both possibilities in the following.
II.1 Neutrino isocurvature from a dark sector
The simplest extension to produce neutrino isocurvature is the setup presented in Fig. 1 (c). A second scalar field $\Psi$ next to the inflaton, which therefore can have uncorrelated quantum perturbations before inflation, later decays to dark radiation which at the time of CMB emission is free-streaming and does not interact with other fluids. This dark sector would necessarily carry uncorrelated isocurvature perturbations.
While neutrino masses are expected to become relevant to LSS surveys in the near future [39, 37, 22], to this day, neutrinos may be modeled as free-streaming radiation, only interacting gravitationally at all times after BBN. Neutrinos and the dark sector are then indistinguishable to CMB and LSS probes. To this end, we define an effective neutrino density and isocurvature perturbation
where we assumed $S_{\gamma\nu}=0$. One may naively think that scenario (c) in the absence of direct interaction (indicated by gray arrows in Fig. 1) corresponds to ${S}_{\gamma\nu_{\mathrm{eff}}}\neq 0\,$ and ${S}_{\gamma\mathrm{DM}}=0$, while (d) corresponds to ${S}_{\mathrm{DM}\,\mathrm{DR}}=0$, which is equivalent to $S_{\gamma\mathrm{DM}}=\rho_{\nu_{\mathrm{eff}}}/\rho_{\mathrm{DR}}\,S_{\gamma\nu_{\mathrm{eff}}}$.
It is obvious that if we allow for direct interactions between the fluids, any ratio of $S_{\gamma\nu_{\mathrm{eff}}}/S_{\gamma\mathrm{DM}}$ may be achieved. For example if we introduce in (c) an interaction that partially thermalizes the DM and DR sector, $S_{\mathrm{DM}\,\mathrm{DR}}$ would be reduced, and depending on the interaction strength, one can interpolate to any value of $S_{\gamma\nu_{\mathrm{eff}}}/S_{\gamma\mathrm{DM}}$ between the ones of example (c) and (d).
In an upcoming paper [28], we will further concentrate on example (c) and show that, contrary to the naive expectation, even in the absence of direct interactions, a non-zero $S_{\gamma\mathrm{DM}},\,S_{\gamma\mathrm{b}}$ can be induced. Using the separate universe approach [51, 46, 44, 9] one can roughly explain the effect as follows: In the presence of dark radiation isocurvature on super horizon scales, the Universe may be viewed as a collection of patches with varying dark radiation density at a given density in the standard model sector (the one originating from $\Phi$ in Fig. 1 (c)). In most models of dark matter production [21] or baryogenesis [25], the yield crucially depends on the ratio between particle physics rates and the Hubble rate. The particle physics rates themselves only depend on the density in the standard model sector (which are usually both given as a function of temperature) and are therefore the same for all patches with the above slicing. The Hubble rate, on the other hand, depends on the energy density in both sectors and therefore varies. This results in varying yields in different patches, a non-vanishing matter isocurvature. Similar effects have been discussed in the context of the curvaton in Refs. [45, 41], and Ref. [4] pointed out that the dark radiation scenario leads to varying element abundances during BBN in the same way.
It should be noted that in the purely gravitationally coupled cases, the amount of induced matter isocurvature is proportional to the relative energy density in the dark sector $\Omega_{\mathrm{DR}}$ at the time of dark matter production and baryogenesis, respectively. The effect might therefore be suppressed if e.g. the dark sector behaves matter-like initially and only later on acts as radiation, resulting in $S_{\gamma\mathrm{DM}}\simeq 0$. This is also the reason why these complications do not appear in case (b) of dark matter isocurvature. Let us also mention that the amount of dark radiation is bounded at the 95% confidence level $\rho_{\mathrm{DR}}/\rho_{\nu}\lesssim 0.1$ from CMB observations [6]. Below we present our bounds in terms of $S_{\gamma\nu_{\mathrm{eff}}}$ such that small dark radiation densities may be equivalently compensated by larger fluctuations $S_{\gamma\mathrm{DR}}$ .
II.2 Neutrino isocurvature from lepton number fluctuations
Let us briefly comment on the scenario without dark radiation present at CMB emission. A new physics sector which decays into highly relativistic neutrinos after they decouple at around $T\approx$1\text{\,}\mathrm{M}\mathrm{e}\mathrm{V}$\$ can be treated in analogy to the previous discussion by identifying the energy in neutrinos resulting from the decay with the free-streaming dark radiation from before.
A lepton number conserving decay at earlier times only leads to an increase of the temperature in the primordial plasma and would not induce neutrino isocurvature. If the decay induces fluctuations in the lepton number to photon ratio, this may still lead to fluctuations in the neutrino energy density. Since this dependence starts at quadratic order only, a sizable background lepton asymmetry is required in order for its fluctuations to induce perturbations in the energy density at linear order [45, 23]. The fluctuations in the energy density again lead to the gravitational effects that induce matter fluctuations. Furthermore, if the large lepton number is linked to the generation of the observed baryon asymmetry, one would expect an imprint of fluctuations on the baryon number as well. In scenarios where the SM plasma reaches temperatures well above the electroweak scale, sphaleron processes equilibrate the lepton and baryon asymmetry, favoring a tiny lepton asymmetry [33]. The generation of the baryon asymmetry would need to be a much weaker residual effect. We therefore reach the same conclusion that neutrino isocurvature is in general accompanied by matter isocurvature. Note that the lepton asymmetry is bounded by CMB and BBN observations [23, 27] paralleling the bound on $\rho_{\mathrm{DR}}$.
II.3 Basis choice for mixed isocurvature
@@
Let us now return to the question whether any specific linear combinations of isocurvature are particularly well motivated. First note that the so called compensated matter perturbation, which is defined by the isocurvature of the total matter content with respect to the total radiation vanishing and corresponds to
@@
is not observable at linear order in the CMB [30, 31]. It is therefore only constrained at amplitudes much larger than the observed adiabatic ones [7, 38, 12]. In this paper we restrict ourselves to magnitudes of isocurvature smaller than the observed CMB fluctuations, $S\lesssim 10^{-5}$. This allows us to recast any fluctuation of baryons and dark matter into an effectively unconstrained compensated matter perturbation and an overall matter fluctuation that may be expressed by a dark matter fluctuation, while choosing $S_{\gamma b}=0$ . This justifies our omission of baryon isocurvature for the most part in the discussion so far and reduces the question of what isocurvature perturbation to consider to a choice of the ratio $S_{\gamma\nu}/S_{\gamma\mathrm{DM}}$ .
@@
Previous works [35, 23, 34, 7, 4, 29] have only considered specific choices of the angle $\varphi$ . The initial condition called dark matter isocurvature corresponds to $\varphi=0$ with our definitions. The one referred to as neutrino isocurvature in the literature is however arranged such that there is no isocurvature between the total radiation and matter. This corresponds to the combination $\varphi~=~\arctan[(1+\rho_{\gamma}/\rho_{\nu})/(1+\rho_{b}/\rho_{\mathrm{DM}})]\approx 1.12$ . Further the normalization is chosen such that the isocurvature fraction $f_{\mathrm{iso}}$ needs to be multiplied by a factor of $1.4$ to compare to our convention. We will refer to this choice of $\varphi$ as compensated neutrino isocurvature in analogy to the compensated matter perturbation. Conversely, we will refer to pure neutrino isocurvature as $S_{\gamma\mathrm{DM}}=0,\,\varphi=\pi/2$ , which results from the scenario shown in Fig. 1 (c) if direct and gravitational interactions between the sectors are negligible.
@@ III.1 Computation of cosmological observables
In order to find the effect of isocurvature on the CMB and other cosmological observables, one needs to solve the evolution equations of the perturbations as the relevant scales enter the horizon. This is commonly done in a fixed gauge. The isocurvature mode in question then corresponds to a set of initial conditions that one has to determine in the gauge of choice. Further, since the evolution equations typically feature singularities as $\tau\rightarrow 0$, one has to analytically expand the solution to finite conformal time $\tau$ before the equations can be solved numerically. In Appendix A we give this expansion in synchronous gauge for $S_{\gamma\mathrm{DM}}=1,\,S_{\gamma\nu}=0$ and $S_{\gamma\mathrm{DM}}=0,\,S_{\gamma\nu}=1$ . The general case may then be obtained by adding the solutions with weights $\cos(\varphi)$ and $\sin(\varphi)$ , respectively. This approach is straightforward to generalize to more exotic scenarios containing e.g. self-interacting instead of free-streaming dark radiation [29].
We use a modified version of the code CLASS [15] to solve the evolution equations of the perturbations and compute all relevant CMB and LSS observables. In Fig. 2 we show the spectrum of lensed temperature correlations in the CMB for a fiducial $\mathrm{\Lambda CDM}$ cosmology using the best fit parameters from Planck [6], but for simplicity taking a flat primordial spectrum, i.e., setting the scalar index $n=1$ . Besides the adiabatic initial conditions that are dominating our Universe, we show four different scenarios with isocurvature instead of the adiabatic mode at the same amplitude.
@@
It becomes apparent that the resulting spectra differ in the position of the acoustic peaks as well as their heights. It is therefore clear that CMB data is able to distinguish between different values of $\varphi$ . In the event of detection, $\varphi$ can therefore be determined and would serve as an invaluable piece of information when trying to disentangle the various cosmologies discussed above. On the other hand, since the signatures differ quiet drastically, it is easy to imagine that an isocurvature signal may be missed when narrowing the search to one $\varphi$ only. The same conclusion is reached in Appendix B, where we consider CMB polarization and LSS.